Hi all, I asked this question recently but havnt been given a responce yet. Any help would be greatly appreciated.
I need to Solve the following system of equations by using Gaussian elimination. I also need to include the row operation used next to each row.
2x - y + 3z = 24
2y - z = 14
2x - 5y + 7z = 6
Thanks, -Nick.

Answer by stanbon(57307)

Solve the following system of equations by using Gaussian elimination. I also need to include the row operation used next to each row.
2x - y + 3z = 24
0x + 2y - z = 14
2x - 5y + 7z = 6
------------------------
Subtract 1st row from 3rd:
2x - y + 3z = 24
0x + 2y - z = 14
0x - 4y + 4z = -18
--------------------------
Add 2 times the 2nd to the 3rd:
2x - y + 3z = 24
0x + 2y - z = 14
0x + 0y + 2z = 10
---------------------
Solve the 3rd equation for "z"
z = 5
----
Substitute into the 2nd Eq and solve for "y":
2y -5 = 14
2y = 19
y = 9.5
------------------
Substitute into the 1st Eq and solve for "x":
2x - 9.5 + 15 = 24
2x + 5.5 = 24
2x = 18.5
x = 9.25
=================================
Cheers,
Stan H.